# An Evaluation of Sakana’s AI Scientist for Autonomous Research: Wishful Thinking or an Emerging Reality Towards ‘Artificial General Research Intelligence’ (AGRI)?

### v1

@@ Abstract.
Abstract. A major milestone toward Artificial General Intelligence (AGI) and even Super Intelligence is AI’s ability to autonomously conduct research—a concept we term Artificial General Research Intelligence (AGRI). If machines could generate hypotheses, design and execute experiments, analyze results, write research papers, and refine knowledge without human intervention, it would transform science. Recently, Sakana.ai introduced the AI Scientist, a system claiming to automate the entire research lifecycle, generating both excitement and skepticism.
We independently evaluated Sakana’s AI Scientist [^1] This study is entirely independent and unaffiliated with Sakana. All statements, evaluations, and findings are based on our own research, analysis, and verification, without any endorsement or influence from Sakana or its representatives. and found it to be a milestone in AI-driven research. While it streamlines some aspects, it falls short of expectations. Literature reviews are weak, nearly half of the experiments failed, and manuscripts sometimes contain hallucinated results. Most notably, users must provide an experimental pipeline as input, limiting the AI Scientist’s autonomy in designing and conducting research. Despite its limitations, the AI Scientist advances research automation. Many reviewers or instructors conducting only a superficial assessment may not recognize its output as AI-generated, instead attributing it to an unmotivated undergraduate. The system produces research papers with minimal human effort and low cost. Our analysis suggests generating a paper costs 666–6 –15 with 3.5 hours of human involvement, making it significantly faster than humans. Compared to AI capabilities from a few years ago, this marks notable progress toward AGRI.
The rise of AI-driven research systems necessitates urgent discussion within Information Retrieval (IR) and broader scientific communities. Enhancing literature retrieval, citation validation, and evaluation benchmarks could improve AI-generated research reliability. We propose concrete steps, including AGRI-specific benchmarks, refined peer review mechanisms, and standardized attribution frameworks. Whether AGRI becomes a stepping stone to AGI depends on how the academic and AI communities shape its development and governance.
@@ 1. Introduction
Notably, just one day the submission of our current manuscript for review to a major IR conference, and hours before uploading it to arXiv, Google announced the AI Co-Scientist [^5] https://blog.google/feed/google-research-ai-co-scientist/, indicating the emerging interest in autonomous research agents, or, as we term it ‘Artificial General Research Intelligence’ (AGRI), i.e. artificial intelligence that has not yet reached the level of general intelligence, but intelligence that is capable of conducting research that is indistinguishable from human research. In this context it is also interesting to note that (Li et al., 2024) submitted a paper on AI-driven idea generation to ICLR 2025, which was rejected after an intense review process involving 45 discussion posts — described by the conference chair as “filled with drama” [^6] https://openreview.net/forum?id=GHJzxPgFa6&noteId=tN3UypVtcw. This case exemplifies the unease among researchers and reviewers who face the prospect of AI encroaching on their roles, raising questions about the future of human-led scientific inquiry.
@@
Notably, the most positive reviewers of the AI Scientist, e.g., (Berman, 2024), have not tested the system themselves but rely solely on Sakana’s information. Meanwhile, users in online forums and GitHub issue trackers report difficulties in setting up the AI Scientist [^8] https://www.reddit.com/r/learnmachinelearning/comments/1fuw2yb/has_anybody_gotten_sakana_ai_scientist_to_work/ or encountering technical issues [^9] https://github.com/SakanaAI/AI-Scientist/issues. Other discussions [^10] https://spectrum.ieee.org/ai-for-science-2 use the AI Scientist’s release as a broader reflection on AI’s role in scientific discovery. Some reviews provide in-depth analyses [^11] https://medium.com/@nimritakoul01/evaluating-the-ai-scientist-63e419e575b8, examining its source code and generated papers. However, to the best of our knowledge, a comprehensive, independent assessment based on direct experimentation has yet to be conducted.
The AI Scientist is a machine learning-driven system, making it highly relevant to the machine learning research community. However, its impact extends beyond ML, with profound implications for the Information Retrieval (IR) community, which motivates us to write this perspective paper. While a recent perspective paper (Zhai, 2024) provided valuable insights into how LLMs are transforming traditional IR tasks, we must also consider a broader question: What happens when LLMs move beyond assisting IR research and begin to conduct it autonomously? This possibility is no longer theoretical — AI systems are increasingly capable of generating research ideas, designing experiments, and even evaluating scientific work.
@@
Our findings reveal a system that, while far from replacing human researchers, demonstrates an ability to automate significant parts of the research process. The AI Scientist does not yet fulfill its promises, struggling with methodological soundness, experimental execution, and literature retrieval. However, these limitations are technical hurdles rather than fundamental barriers — challenges that will likely be addressed as AI-driven research systems evolve.
@@ 2.1. Setup and Installation
All coding and experiment execution were carried out by the third author, a third-year computer science Bachelor’s student with basic machine learning knowledge but strong Python proficiency. A typical user of the AI Scientist is likely to have greater expertise, such as that of a trained researcher or machine learning engineer. Thus, our setup time and usability assessments should be considered lower bounds. We have made our experimental codebase that replicates our experimentation in this paper available on GitHub: https://code.isg.beel.org/ISG-Siegen/the-ai-scientist-reproduced. The repository includes a Singularity container definition file to facilitate full reproducibility and modifications in our experiments.
@@
We first tested and prototyped the AI Scientist on a consumer laptop [^13] HUAWEI MateBook D 16, Windows 11, Intel(R) Core(TM) i5-12450H with integrated graphics, 16GB RAM (3733 MT/s). The main experiments were then conducted on our university’s computing cluster [^14] Single node with 2 AMD EPYC 7452 CPUs (32 cores, 2.35-3.35 GHz, 128 MB cache) and 256 GB DDR4 RAM (3200 MHz), running Rocky Linux 8.8. Although the AI Scientist provides an experimental pipeline optimized for deep learning with PyTorch and an NVIDIA GPU (CUDA), our experiments used a simple recommendation algorithm, FunkSVD, that runs on a CPU.
@@ 2.2. Preparing for the experiments
We initially assumed the AI Scientist could autonomously conduct research based solely on a prompt. However, it requires a user-defined “template,” which significantly limits its autonomy. Such a template consist of several elements, described in the following.
@@ 1. Goal and Research Direction
The research goal is specified via a prompt file (prompts.json). For our experiments, we focused on Green Recommender Systems, a research topic which aims to reduce the carbon footprint of recommender systems and its associated machine learning algorithms. We gave the prompt as:  [^15] All prompts are shortened and rephrased for brevity; the originals are available in our GitHub repository.:
@@
2 "task_description": "The attached code trains and evaluates a functional svd matrix factorization with stochastic gradient descent for recommender systems. Your goal is to find novel ways to optimize energy-efficiency. "
@@ 2. Experimental Pipeline
The AI Scientist requires a Python-based experimental pipeline, consisting of an experiment.py, which defines dataset usage and model training as well as a plot.py, which handles result visualization. We implemented FunkSVD for collaborative filtering, training and evaluating it on an 80/10/10 split on MovieLens-100k (RMSE metric), with results visualized via line charts. This minimalist setup runs on a CPU, but larger datasets and complex models would likely increase compute costs.
@@ 3. Seed Ideas
    Adaptive Learning Rates for SGD. Aims to accelerate convergence, reducing computational cost and energy use. Though not novel, it demonstrates our intended research direction.
@@
    E-Fold Cross-Validation. An alternative to k-fold cross-validation that dynamically selects an optimal $e$ to balance computational efficiency and reliability (Mahlich et al., 2024; Beel et al., 2024; Baumgart et al., 2024).
@@ 2.3. Idea Generation
Once provided with seed ideas and an experimental template, the AI Scientist generated the following ten research ideas:
@@ 10. Hybrid matrix factorization
Despite mimicking a literature review, this approach is unreliable. In our experiments, the AI Scientist classified all 10 generated ideas and both seed ideas as novel [^16] In prior tests, the AI Scientist occasionally flagged some ideas as non-novel, showing that it does not always default to a “novel” classification., despite some being well-documented. For instance, micro-batching for SGD (Idea 7) is a known technique (Jain et al., 2018), and hybrid matrix factorization, though reasonable, is neither novel nor directly related to energy efficiency. This misclassification likely results from the AI Scientist’s reliance on keyword matching rather than any deeper synthesis of research. While some ideas may be novel within recommender systems, many simply apply existing techniques (e.g., pruning) to matrix factorization, offering incremental rather than groundbreaking improvements.
@@ 2.4. Conducting Experiments
Our code analysis suggests minimal changes per iteration. Our template had 255 lines (6,260 characters), with Aider adding, on average, 529 characters (+8%) in the first iteration, followed by 118, 83, 66, and 21 characters in subsequent iterations (see Table 1).
@@ 10 Hybrid MF Failed to run
The AI Scientist struggled with both implementation and methodological correctness. It failed to execute five of twelve ideas due to unresolved coding errors, often cycling through flawed versions of the same code. Even when producing executable code, outputs were frequently unreliable. For instance, in our e-fold cross-validation experiment, the AI Scientist intended to test $e=\{2,3,4,5\}$ but erroneously kept $e$ fixed at 2, invalidating results. It also failed to re-run the baseline, making its novel approach appear superior – an impossible outcome because e-fold cross-validation by design cannot lead to better performance but only energy savings (with, ideally, comparable performance) (Mahlich et al., 2024; Beel et al., 2024; Baumgart et al., 2024). The apparent performance gain may have been caused by statistical noise or incorrect baseline execution.
@@ 2.5. Manuscripts
The AI Scientist generated manuscripts for all seven successfully conducted experiments, with lengths ranging from six to eight pages (median: seven). Most manuscripts (6 of 7) included tables or charts, though quality varied. References were scarce, with a median of five citations per paper (range: 2–9), mostly outdated; only five of the total 34 citations were from 2020 or later. Foundational works, such as Goodfellow et al.’s Deep Learning textbook, were frequently cited, suggesting weak retrieval of relevant literature.
Structural issues were common: four manuscripts contained missing or misplaced figures, incomplete sections (e.g., “Conclusions Here” placeholder), duplicates, or repeated figures. Related work sections were particularly poor, often with irrelevant citations. In our e-fold cross-validation experiment, the AI Scientist failed to cite existing papers on the topic that used exactly the same terminology (e-fold cross validation), despite their availability on Semantic Scholar [^18] https://www.semanticscholar.org/search?q=%22e-fold%20cross%20validation.
Results sections were often inadequate or misleading. Metrics were presented in plain text or simple charts, with no confidence intervals or p-values. Four of seven manuscripts contained incorrect or hallucinated numerical results, with discrepancies in hyperparameters and performance metrics. In one case, an energy-efficient machine learning experiment claimed an RMSE improvement while reducing training time from 116 to 115 seconds but increasing memory usage, failing to justify how this aligned with energy efficiency. Similar logical inconsistencies appeared in other experiments.
Given these flaws, these manuscripts would likely be rejected at reputable conferences or journals. Even at lower-tier venues, the lack of proper citations and unreliable results would be problematic. However, an undergraduate submitting such work for a course project might still pass if the instructor does not scrutinize the code or experiments. The AI Scientist’s output resembles that of a student rushing to complete an assignment – formally structured but careless in execution.
@@ 2.6. Review Functionality
The AI Scientist highlighted valid methodological concerns, such as weak theoretical justification and potential biases, but overlooked critical flaws, including redundant text, formatting errors, missing sections, and flawed experimental results. While human reviews vary in quality – from superficial comments to detailed critiques – the AI Scientist’s reviews were consistently structured, though lacking depth. Despite this, they appeared plausible enough that editors and conference chairs might not immediately recognize them as AI-generated. While less rigorous than strong human reviews, they were still more substantive than the brief, low-effort reviews that sometimes occur in the conference reviewing process.
@@
A key challenge is determining how to evaluate AI-generated reviews, given the inconsistencies of human peer review. Should human judgment be the benchmark? Peer review is known for subjectivity, reviewer disagreements, and biases. An AI system that diverges from human reviewers may still highlight valid weaknesses or apply overly rigid criteria.
This raises a fundamental question: should AI peer review aim to mimic human reviewers or provide a more systematic alternative? If the former, the AI Scientist falls short. If the latter, human evaluation alone may be an insufficient benchmark. A better assessment would compare AI-generated reviews against expert evaluations that measure consistency and validity.

### v2

# Evaluating Sakana’s AI Scientist for Autonomous Research: Wishful Thinking or an Emerging Reality Towards ‘Artificial Research Intelligence’ (ARI)?

@@ Abstract.
Abstract. Recently, Sakana.ai introduced the AI Scientist, a system claiming to automate the entire research lifecycle and conduct research autonomously, a concept we term Artificial Research Intelligence (ARI). Achieving ARI would be a major milestone toward Artificial General Intelligence (AGI) and a prerequisite to achieving Super Intelligence. The AI Scientist received much attention in the academic and broader AI community. A thorough evaluation of the AI Scientist, however, had not yet been conducted.
We evaluated the AI Scientist and found several critical shortcomings. The system’s literature review process is inadequate, relying on simplistic keyword searches rather than profound synthesis, which leads to poor novelty assessments. In our experiments, several generated research ideas were incorrectly classified as novel, including well-established concepts such as micro-batching for stochastic gradient descent (SGD). The AI Scientist also lacks robustness in experiment execution—five out of twelve proposed experiments (42%) failed due to coding errors, and those that did run often produced logically flawed or misleading results. In one case, an experiment designed to optimize energy efficiency reported improvements in accuracy while consuming more computational resources, contradicting its stated goal. Furthermore, the system modifies experimental code minimally, with each iteration adding only 8% more characters on average, suggesting limited adaptability. The generated manuscripts were poorly substantiated, with a median of just five citations per paper—most of which were outdated (only five out of 34 citations were from 2020 or later). Structural errors were frequent, including missing figures, repeated sections, and placeholder text such as “Conclusions Here”. Hallucinated numerical results were contained in several manuscripts, undermining the reliability of its outputs.
Despite its limitations, the AI Scientist represents a significant leap forward in research automation. It produces complete research manuscripts with minimal human intervention, challenging conventional expectations of AI-generated scientific work. Many reviewers or university instructors conducting only a superficial assessment may struggle to distinguish its output from that of human researchers, demonstrating how far AI has progressed in mimicking academic writing and structuring scientific arguments. While the quality of its manuscripts currently aligns with that of an unmotivated undergraduate student rushing to meet a deadline, this level of autonomy in research generation is remarkable. More strikingly, it achieves this at an unprecedented speed and cost efficiency—our analysis indicates that generating a full research paper costs only \$6–\$15, with just 3.5 hours of human involvement. This is significantly faster than traditional human researchers. Given that AI research automation was nearly nonexistent just a few years ago, the AI Scientist marks a substantial milestone toward Artificial Research Intelligence (ARI), signalling the acceleration of AI-driven scientific discovery.
The AI Scientist also illustrates the urgent need for a discussion within the Information Retrieval (IR) and broader scientific communities. Whether and when ARI becomes a reality depends on how the academic and AI communities shape its development and governance. We propose concrete steps, including pilot projects and competitions, and standardized attribution frameworks such as research logs and markup languages.
@@ 1. Introduction
Notably, just one day before the submission of our current manuscript for review to a major IR conference, and hours before uploading it to arXiv, Google announced the AI Co-Scientist [^4] https://blog.google/feed/google-research-ai-co-scientist/, indicating the emerging interest in autonomous research agents, or, as we term it ‘Artificial Research Intelligence’ (ARI), i.e. artificial intelligence that has not yet reached the level of general intelligence, but intelligence that is capable of conducting research that is indistinguishable from human research. In this context it is also interesting to note that (Li et al., 2024) submitted a paper on AI-driven idea generation to ICLR 2025, which was rejected after an intense review process involving 45 discussion posts — described by the conference chair as “filled with drama” [^5] https://openreview.net/forum?id=GHJzxPgFa6&noteId=tN3UypVtcw. This case exemplifies the unease among researchers and reviewers who face the prospect of AI encroaching on their roles, raising questions about the future of human-led scientific inquiry.
@@
Notably, the most positive reviewers of the AI Scientist, e.g., (Berman, 2024), have not tested the system themselves but rely solely on Sakana’s information. Meanwhile, users in online forums and the GitHub issue trackers report difficulties in setting up the AI Scientist [^7] https://www.reddit.com/r/learnmachinelearning/comments/1fuw2yb/has_anybody_gotten_sakana_ai_scientist_to_work/ or encountering technical issues [^8] https://github.com/SakanaAI/AI-Scientist/issues. Other discussions [^9] https://spectrum.ieee.org/ai-for-science-2 use the AI Scientist’s release as a broader reflection on AI’s role in scientific discovery. Some reviews provide in-depth analyses [^10] https://medium.com/@nimritakoul01/evaluating-the-ai-scientist-63e419e575b8, examining its source code and generated papers. However, to the best of our knowledge, a comprehensive, independent assessment based on direct experimentation has yet to be conducted.
The AI Scientist is a machine learning driven system, making it highly relevant to the machine learning research community. However, its impact extends beyond machine learning, with profound implications for the Information Retrieval (IR) community, which motivates us to write this paper. While a recent SIGIR perspective paper (Zhai, 2024) provided valuable insights into how LLMs are transforming traditional IR tasks, we must also consider a broader question: What happens when LLMs move beyond assisting IR research and begin to conduct it autonomously? This possibility is no longer theoretical — AI systems are increasingly capable of generating research ideas, designing experiments, and even evaluating scientific work.
@@
Our findings reveal a system that, while far from replacing human researchers, demonstrates the potential to automate significant parts of the research process. The AI Scientist does not yet fulfill its promises, struggling with methodological soundness, experimental execution, and literature retrieval. However, these limitations are technical hurdles rather than fundamental barriers — challenges that will likely be addressed as AI-driven research systems evolve.
@@ 2.1. Setup and Installation
All setup, coding and experiment execution were carried out by the third author of this manuscript, a third-year computer science Bachelor’s student with basic machine learning knowledge but strong Python proficiency. A typical user of the AI Scientist is likely to have greater expertise, such as that of a trained researcher or machine learning engineer. Thus, our setup time and usability assessments should be considered lower bounds. We have made our experimental codebase that replicates our experimentation in this paper available on GitHub: https://code.isg.beel.org/ISG-Siegen/the-ai-scientist-reproduced. The repository includes a Singularity container definition file to facilitate full reproducibility and modifications in our experiments.
@@
We first tested and prototyped the AI Scientist on a consumer laptop [^12] HUAWEI MateBook D 16, Windows 11, Intel Core i5-12450H with integrated graphics, 16GB RAM (3733 MT/s). The main experiments were then conducted on our university’s computing cluster [^13] Single node with 2 AMD EPYC 7452 CPUs (32 cores, 2.35-3.35 GHz, 128 MB cache) and 256 GB DDR4 RAM (3200 MHz), running Rocky Linux 8.8. Although the AI Scientist provides an experimental pipeline optimized for deep learning with PyTorch and an NVIDIA GPU (CUDA), our experiments used a simple recommendation algorithm, FunkSVD, that runs on a CPU.
@@ 2.2. Preparing for the experiments
We initially assumed the AI Scientist could autonomously conduct research based solely on a prompt. However, it requires a user-defined “template,” which significantly limits the autonomy of the AI Scientist. Such a template consist of several elements, described in the following.
@@ 1. Goal and Research Direction
The research goal is specified via a prompt file (prompts.json). For our experiments, we focused on Green Recommender Systems (Beel et al., 2024a), a research topic which aims to reduce the carbon footprint of recommender systems and its associated machine learning algorithms. We gave the prompt as:  [^14] All prompts are shortened and rephrased for brevity; the originals are available in our GitHub repository.:
@@
2 "task_description": "The attached code trains and evaluates the FunkSVD algorithm with stochastic gradient descent for recommender systems. Your goal is to find novel ways to optimize energy-efficiency. "
@@ 2. Experimental Pipeline
The AI Scientist requires a Python-based experimental pipeline, consisting of an experiment.py, which defines dataset usage and model training as well as a plot.py, which handles result visualization. We implemented FunkSVD for collaborative filtering, training and evaluating it on an 80/10/10 split on MovieLens-100k (RMSE metric), with results visualized via line charts. This minimalist setup runs on a CPU, but larger datasets and complex models would likely increase compute costs. It is important to note that the AI Scientist requires this pipeline in a special format. Users cannot simply input any Python code but would have to adjust their existing code to be able to be used by the AI Scientist.
@@ 3. Seed Ideas
    Adaptive Learning Rates for SGD. Aims to accelerate convergence, reducing computational cost and energy use. Though not novel, this ideas demonstrates our intended research direction.
@@
    E-Fold Cross-Validation. An alternative to k-fold cross-validation that dynamically selects an optimal $e$ to balance computational efficiency and reliability. This idea is also not entirely novel, but was proposed only recently(Mahlich et al., 2024; Beel et al., 2024b; Baumgart et al., 2024).
@@ 2.3. Idea Generation
Once provided with required input, the AI Scientist generated the following ten research ideas:
@@ 10. Hybrid matrix factorization
Despite mimicking a literature review, this approach is unreliable. In our experiments, the AI Scientist classified all 10 generated ideas and both seed ideas as novel [^15] In prior tests, the AI Scientist occasionally flagged some ideas as non-novel, showing that it does not always default to a “novel” classification., despite some being well-documented. For instance, micro-batching for SGD (Idea 7) is a known technique (Jain et al., 2018), as are adaptive learning rates (seed idea), hybrid matrix factorization (idea 10) and e-fold cross validation (seed idea). This misclassification likely results from the AI Scientist’s reliance on keyword matching rather than any deeper synthesis of research. While some ideas may be novel within recommender systems, many simply apply existing techniques (e.g., pruning) to matrix factorization, offering incremental rather than groundbreaking improvements.
@@ 2.4. Conducting Experiments
Our code analysis suggests minimal changes per iteration. Our template had 6,260 characters of code (255 lines), with Aider adding, on average, 529 characters (+8%) in the first iteration, followed by 118, 83, 66, and 21 characters in subsequent iterations (see Table 1).
@@ 10 Hybrid MF Failed to run
The AI Scientist struggled with both implementation and methodological correctness. It failed to execute five of twelve ideas due to unresolved coding errors, often cycling through flawed versions of the same code. Even when producing executable code, outputs were frequently unreliable. For instance, in our e-fold cross-validation experiment, the AI Scientist intended to test $e=\{2,3,4,5\}$ but erroneously kept $e$ fixed at 2, invalidating results. It also failed to re-run the baseline (k-fold), making e-fold cross validation appear superior – an impossible outcome because e-fold cross-validation by design cannot lead to better performance than k-fold cross validation but only energy savings (with, ideally, comparable performance) (Mahlich et al., 2024; Beel et al., 2024b; Baumgart et al., 2024). There are many more similar examples that are available in our GitHub repository.
@@ 2.5. Manuscripts
The AI Scientist generated manuscripts for all seven successfully conducted experiments, with lengths ranging from six to eight pages (median: seven). Most manuscripts (6 of 7) included tables or charts, though quality varied. References were scarce, with a median of five references per paper (range: 2–9), mostly outdated; only five (14.7%) of the total 34 references were from 2020 or later. Foundational works, such as Goodfellow et al.’s Deep Learning textbook, were frequently cited, suggesting weak retrieval of relevant literature.
Structural issues were common: four manuscripts (57%)contained missing or misplaced figures, incomplete sections (e.g., “Conclusions Here” placeholder), duplicates, or repeated figures. Related work sections were particularly poor in all manuscripts, often with irrelevant citations. In our e-fold cross-validation experiment, the AI Scientist failed to cite existing papers on the topic that used exactly the same terminology (e-fold cross validation), despite their availability on Semantic Scholar [^17] https://www.semanticscholar.org/search?q=%22e-fold%20cross%20validation.
Results sections were often inadequate or misleading. Metrics were presented in plain text or simple charts, with no confidence intervals or p-values. Four of seven manuscripts (57%) contained incorrect or hallucinated numerical results, with discrepancies in hyperparameters and performance metrics. In one case, the AI Scientist claimed to have successfully developed an energy-efficient machine learning algorithm, supposedly evidenced by better RMSE and reduction of training time from 116 to 115 seconds – but increased memory usage. How an algorithm that is optimized for energy usage (or training time) may lead to better RMSE was not discussed (one would expect the opposite). Neither was discussed if the one-second-improvement (or 1%), in training time was statistically (or practically) significant; nor was the implication discussed that memory usage increased. Similar logical inconsistencies appeared in other experiments.
Given these flaws, these manuscripts would likely be rejected at reputable conferences or journals. Even at lower-tier venues, the lack of proper citations and unreliable results would be problematic. However, an undergraduate submitting such work for a course project might still pass if the instructor does not scrutinize the code or experiments. Therefore, in our opinion, the AI Scientist’s output resembles that of a student rushing to complete an assignment – formally structured but careless in execution.
@@ 2.6. Review Functionality
The AI Scientist highlighted more valid concerns, such as weak theoretical justification and potential biases, but overlooked critical flaws, including redundant text, formatting errors, missing sections, and flawed experimental results. Most notably, the reviewer agent did not identify any of the truly serious issues that we highlighted in the sections above.
While human reviews vary in quality – from superficial comments to detailed critiques – the AI Scientist’s reviews were consistently structured, though lacking depth. Despite this, they appeared plausible enough that editors and conference chairs might not immediately recognize them as AI-generated. While less rigorous than strong human reviews, they were still more substantive than the brief, low-effort reviews that sometimes occur in the conference reviewing process.
@@
Our evaluation of the AI Scientist reviewer agent, illustrates a challenge: how to evaluate AI-generated reviews, given the inconsistencies of human peer review? Should human judgment be the benchmark? Peer review is known for subjectivity, reviewer disagreements, and biases. An AI system that diverges from human reviewers may still highlight valid weaknesses or apply overly rigid criteria.
This raises another fundamental question: should AI peer review aim to mimic human reviewers or provide a more systematic alternative? If the former, the AI Scientist falls short. If the latter, human evaluation alone may be an insufficient benchmark. A better assessment would compare AI-generated reviews against expert evaluations that measure consistency and validity.